# Setting $\varepsilon$ is not the Issue in Differential Privacy

### v1

@@ 4.2 Rejecting if not smallest epsilon
Rejecting papers solely for not achieving state-of-the-art (SOTA) results is widely recognized as a bad reviewing practice, as it replaces the difficult task of assessing a work’s value to the community with a simple numerical comparison. This approach fails to consider whether worse performance on some metrics arises from intrinsic limitations or from trade-offs worth exploring. It can limit the publication of research directions that are not yet competitive with mainstream methods but have high potential, ultimately harming research diversity bender2021dangers; raji2021ai; Church_Kordoni_2022.
@@
Finally, as already developed in Section 3 for the emoji computation, moderately large $\varepsilon$ can still correspond to meaningful protection as the number of possible outputs increases. Intuitively, as the dimension grows, the probability space becomes hard to constrain but also hard to explore simultaneously, so crafting one of the worst measurable sets $\mathcal{S}$ and estimating its probability becomes harder as well. If one is interested in protecting against reconstruction rather than membership inference, current evidence suggests that an attacker cannot reconstruct the dataset as long as the privacy budget remains small with respect to the dimension inproceedings. In other words, dismissing algorithms with a large constant privacy budget may limit our understanding of more complex tasks and models, while no alternative privacy protections with theoretical guarantees currently exist for this scenario.
@@ 5.1 Promoting more "explicit" methods
Protecting data is intrinsically hard, and many intuitive methods are flawed, as already discussed in Section˜2. However, this is not always known by newcomers or practitioners who are primarily looking for a practical solution. The idea that there is a hyperparameter $\varepsilon$ that must be set to some elusive “magic” value, which no algorithm can currently determine for you, can feel intimidating. In contrast, other techniques such as similarity-based privacy metrics may initially seem more appealing. However, their supposed advantages are in fact non-existent, as they do not actually protect the data Emiliano, and can even lead to a false sense of security, or worse, to additional leakage desfontainesblog20240606.
These methods are often claimed to be easier to explain, but this advantage does not hold under close scrutiny. In fact, differential privacy now benefits from a large body of independent empirical attack studies conducted by unrelated research teams, which provides a more balanced and credible assessment than attacks designed by the same teams who proposed the defense quantifyeps; tablereidentification; nanayakkara2023epsilon; epswithattributeattacks. This independent validation helps build a more objective understanding of the strengths and limitations. This point is unlikely to be controversial within the privacy community, which is well aware of the lack of scientific foundation behind these alternative approaches. However, people outside the field may lack the broader context, especially when differential privacy is portrayed as impractical due to overly stringent expectations about privacy budget selection. In particular, there is a troubling asymmetry in the burden of proof: it is sometimes implicitly assumed that the relevance of differential privacy has already been disproven, while alternative methods are not held to the same standard. This framing shifts the responsibility onto privacy researchers to find successful attacks against these methods, rather than requiring proponents to establish their robustness in the first place.
Placing the burden of proof on privacy researchers is impractical, as each flawed method may require a different attack strategy, and there is little incentive to conduct such work. These results are often of limited scientific interest, and even when a method is clearly shown to be broken, this rarely deters its continued promotion – illustrated, for example, by the case of InstaHide, which continue to promote its privacy guarantees depsite the possible full reconstruction of the data pmlr-v119-huang20i; Carlini2020AnAO. We should reverse the burden of proof in discussions aimed at general audiences and make it clear that any claim of data privacy that does not rely on differential privacy must be explicitly justified.
@@ 5.2 Deploying privacy-washing implementations
Another risk of popularizing the idea that privacy budgets are hard to set is that it can make weakly private algorithms appear more protective than they actually are. The danger of presenting privacy as a selling point despite offering little real guarantee was highlighted in reactions to Apple’s first deployment of differential privacy. In that case, a budget of $16$ was allowed for each individual day, which meant that after just a few days, the cumulative budget exceeded what could be considered a meaningful privacy guarantee tang2017privacylossapplesimplementation. It is important to note that the advertised privacy budget was a small value – $1$, which was later shown to degrade significantly due to composition. This example illustrates how confusion around privacy budgets can enable misleading claims and poor implementations to go unchecked.
@@ 5.3 Delaying until DP is solved
Finally, emphasizing current limitations of differential privacy too heavily may lead to postponing adoption, as "no one really knows how to protect data yet." But it is unrealistic to expect all open problems in differential privacy to be solved, especially since every new advance in machine learning tends to raise new privacy challenges. Differential privacy has already been shown to be effective in various real-world deployments.10.1145/3294052.3322188; Erlingsson2014RAPPORRA. The clear and urgent message from the community should be to encourage broader adoption, rather than to focus primarily on limitations.
@@ 6 Alternative views
One might argue that $(\varepsilon,\delta)$-differential privacy is not precise enough compared to other metrics such as Rényi DP or, more recently, Gaussian DP as many mechanisms are not tightly analyzed under standard DP. We clarify that our position is not to discourage the use of tighter accounting methods. Rather, we advocate using them internally and translating the final results into $(\varepsilon,\delta)$-DP for ease of comparison. As we have argued that the privacy budget should primarily be interpreted in terms of its order of magnitude, having a slightly loose conversion to DP is not a serious concern. We see the development of metrics that convert to DP as complementary, not opposed, to the DP framework. These alternative metrics share the same foundational principles, and if they can be converted to $(\varepsilon,\delta)$-DP, they should be considered equivalent in spirit. For instace, Rényi differential privacy (RDP), Gaussian differential privacy (GDP), and $f$-differential privacy ($f$-DP) all admit explicit conversions to $(\varepsilon,\delta)$-differential privacy. Concretely, for any target $\delta>0$, $(\alpha,\varepsilon)$-RDP implies $(\varepsilon+\frac{\log(1/\delta)}{\alpha-1},\delta)$-DP renyi, $\mu$-GDP implies $(\mu\sqrt{2\log(1/\delta)},\delta)$-DP DongRothmanGong2022GDP, and any $f$-DP guarantee induces a corresponding $(\varepsilon,\delta)$-DP. These conversion results show that $(\varepsilon,\delta)$-DP can serve as a common reporting format for a wide range of differential privacy definitions.
@@
This argument is sometimes framed as a strawman, suggesting that the need to set a privacy budget is just another sign that privacy is fundamentally incompatible with machine learning. More broadly, it reflects the belief that trustworthy constraints hinder ML progress, a stance that can appear openly adversarial. However, we believe this view cannot be ignored; for instance, similar arguments have recently surfaced in calls for less alignment in large language models. A common counterargument is that differential privacy enables the use of data that would otherwise be inaccessible. While this may be pragmatically true, it is a dangerous position if it treats privacy merely as a means to achieve wider data collection, rather than as a value in its own right. Such reasoning risks legitimizing problematic use cases under the banner of privacy. Instead, we prefer to refer to philosophical works that highlights the intrinsic importance of privacy for individuals and its deep interconnections with other fundamental rights westin1967privacy; solove2008understanding. Data leakage is often irreversible, and the absence of strong privacy protections poses a serious societal threat.
@@ Acknowledgments
This research was funded in whole by the Austrian Science Fund (FWF) 10.55776/COE12. I thank Alberto Naibo for the discussion and his invitation to "Logique, droit, IA : penser les algorithmes" colloquium. I thank Simone Bombari for the discussions.

### v2

@@ 4.2 Rejecting if not smallest epsilon
Rejecting papers solely for not achieving state-of-the-art (SOTA) results is widely recognized as a bad reviewing practice, as it replaces the difficult task of assessing a work’s value to the community with a simple numerical comparison. This approach fails to consider whether worse performance on some metrics arises from intrinsic limitations or from trade-offs worth exploring. It can limit the publication of research directions that are not yet competitive with mainstream methods but have high potential, ultimately harming research diversity Bender et al. (2021); Raji et al. (2021); Church and Kordoni (2022).
@@
Finally, as already developed in Section 3 for the emoji computation, moderately large $\varepsilon$ can still correspond to meaningful protection as the number of possible outputs increases. Intuitively, as the dimension grows, the probability space becomes hard to constrain but also hard to explore simultaneously, so crafting one of the worst measurable sets $\mathcal{S}$ and estimating its probability becomes harder as well. If one is interested in protecting against reconstruction rather than membership inference, current evidence suggests that an attacker cannot reconstruct the dataset as long as the privacy budget remains small with respect to the dimension Balle et al. (2022). In other words, dismissing algorithms with a large constant privacy budget may limit our understanding of more complex tasks and models, while no alternative privacy protections with theoretical guarantees currently exist for this scenario. On the contrary, it would be interesting to theoretically understand why large $\varepsilon$ provide good empirical defenses in various scenarios Aerni et al. (2024); Nasr et al. (2021); Bhowmick et al. (2019).
@@ 5.1 Promoting more "explicit" methods
Protecting data is intrinsically hard, and many intuitive methods are flawed, as already discussed in Section˜2. However, this is not always known by newcomers or practitioners who are primarily looking for a practical solution. The idea that there is a hyperparameter $\varepsilon$ that must be set to some elusive “magic” value, which no algorithm can currently determine for you, can feel intimidating. In contrast, other techniques such as similarity-based privacy metrics may initially seem more appealing. However, their supposed advantages are in fact non-existent, as they do not actually protect the data Ganev and De Cristofaro (2025), and can even lead to a false sense of security, or worse, to additional leakage Desfontaines (2024).
These methods are often claimed to be easier to explain, but this advantage does not hold under close scrutiny. In fact, differential privacy now benefits from a large body of independent empirical attack studies conducted by unrelated research teams, which provides a more balanced and credible assessment than attacks designed by the same teams who proposed the defense Lee and Clifton (2011); Jayaraman and Evans (2019); Nanayakkara et al. (2023); Pankova and Laud (2022). This independent validation helps build a more objective understanding of the strengths and limitations. This point is unlikely to be controversial within the privacy community, which is well aware of the lack of scientific foundation behind these alternative approaches. However, people outside the field may lack the broader context, especially when differential privacy is portrayed as impractical due to overly stringent expectations about privacy budget selection. In particular, there is a troubling asymmetry in the burden of proof: it is sometimes implicitly assumed that the relevance of differential privacy has already been disproven, while alternative methods are not held to the same standard. This framing shifts the responsibility onto privacy researchers to find successful attacks against these methods, rather than requiring proponents to establish their robustness in the first place.
Placing the burden of proof on privacy researchers is impractical, as each flawed method may require a different attack strategy, and there is little incentive to conduct such work. These results are often of limited scientific interest, and even when a method is clearly shown to be broken, this rarely deters its continued promotion – illustrated, for example, by the case of InstaHide, which continue to promote its privacy guarantees depsite the possible full reconstruction of the data Huang et al. (2020); Carlini et al. (2020). We should reverse the burden of proof in discussions aimed at general audiences and make it clear that any claim of data privacy that does not rely on differential privacy must be explicitly justified.
@@ 5.2 Deploying privacy-washing implementations
Another risk of popularizing the idea that privacy budgets are hard to set is that it can make weakly private algorithms appear more protective than they actually are. The danger of presenting privacy as a selling point despite offering little real guarantee was highlighted in reactions to Apple’s first deployment of differential privacy. In that case, a budget of $16$ was allowed for each individual day, which meant that after just a few days, the cumulative budget exceeded what could be considered a meaningful privacy guarantee Tang et al. (2017). It is important to note that the advertised privacy budget was a small value – $1$, which was later shown to degrade significantly due to composition. This example illustrates how confusion around privacy budgets can enable misleading claims and poor implementations to go unchecked.
@@ 5.3 Delaying until DP is solved
Finally, emphasizing current limitations of differential privacy too heavily may lead to postponing adoption, as "no one really knows how to protect data yet." But it is unrealistic to expect all open problems in differential privacy to be solved, especially since every new advance in machine learning tends to raise new privacy challenges. Differential privacy has already been shown to be effective in various real-world deployments.Dwork (2019); Erlingsson et al. (2014). The clear and urgent message from the community should be to encourage broader adoption, rather than to focus primarily on limitations.
@@ 6 Alternative views
One might argue that $(\varepsilon,\delta)$-differential privacy is not precise enough compared to other metrics such as Rényi DP or, more recently, Gaussian DP as many mechanisms are not tightly analyzed under standard DP. We clarify that our position is not to discourage the use of tighter accounting methods. Rather, we advocate using them internally and translating the final results into $(\varepsilon,\delta)$-DP for ease of comparison. As we have argued that the privacy budget should primarily be interpreted in terms of its order of magnitude, having a slightly loose conversion to DP is not a serious concern. We see the development of metrics that convert to DP as complementary, not opposed, to the DP framework. These alternative metrics share the same foundational principles, and if they can be converted to $(\varepsilon,\delta)$-DP, they should be considered equivalent in spirit. For instace, Rényi differential privacy (RDP), Gaussian differential privacy (GDP), and $f$-differential privacy ($f$-DP) all admit explicit conversions to $(\varepsilon,\delta)$-differential privacy. Concretely, for any target $\delta>0$, $(\alpha,\varepsilon)$-RDP implies $(\varepsilon+\frac{\log(1/\delta)}{\alpha-1},\delta)$-DP Mironov (2017), $\mu$-GDP implies $(\mu\sqrt{2\log(1/\delta)},\delta)$-DP Dong et al. (2022), and any $f$-DP guarantee induces a corresponding $(\varepsilon,\delta)$-DP. These conversion results show that $(\varepsilon,\delta)$-DP can serve as a common reporting format for a wide range of differential privacy definitions.
@@
This argument is sometimes framed as a strawman, suggesting that the need to set a privacy budget is just another sign that privacy is fundamentally incompatible with machine learning. More broadly, it reflects the belief that trustworthy constraints hinder ML progress, a stance that can appear openly adversarial. However, we believe this view cannot be ignored; for instance, similar arguments have recently surfaced in calls for less alignment in large language models. A common counterargument is that differential privacy enables the use of data that would otherwise be inaccessible. While this may be pragmatically true, it is a dangerous position if it treats privacy merely as a means to achieve wider data collection, rather than as a value in its own right. Such reasoning risks legitimizing problematic use cases under the banner of privacy. Instead, we prefer to refer to philosophical works that highlights the intrinsic importance of privacy for individuals and its deep interconnections with other fundamental rights Westin (1967); Solove (2008). Data leakage is often irreversible, and the absence of strong privacy protections poses a serious societal threat.
@@ Acknowledgments
This research was funded in whole by the Austrian Science Fund (FWF) 10.55776/COE12. I thank Alberto Naibo for the discussion and his invitation to "Logique, droit, IA : penser les algorithmes" colloquium. I thank Simone Bombari for the discussions. I thank EurIPS for the active poster session.